OpenStudy (anonymous):

simplify |n^2+1/n^2-1..-1|

OpenStudy (anonymous):

is that absolute value? like this: \(\large |\frac{n^2+1}{n^2-1}-1| \) ????

OpenStudy (anonymous):

yes sorry thats the best i can display it with the keyboard

OpenStudy (anonymous):

ok.... so let's just change that "1" to \(\large \frac{n^2-1}{n^2-1} \) and subtract those fractions... can you tell me what you get?

OpenStudy (anonymous):

change this "1" : \(\large |\frac{n^2+1}{n^2-1}-\color {red}1| \)

OpenStudy (anonymous):

i get 2/n^2-1

OpenStudy (anonymous):

uh... what's the two dots there right before the last "-1" in the original expression?

OpenStudy (anonymous):

i was just trying to separate the -1 in the expression. i have a couple more of these and it seems as though 1 method does not work for the other for example |n-1/3n-1/3| can you help with that

OpenStudy (anonymous):

if they don't mean anything.... then that's what i got.... so the answer is \(\large |\frac{2}{n^2-1}| \) where \(\large n \ne \pm1 \)

OpenStudy (anonymous):

that one is easier than the last one i know

OpenStudy (anonymous):

can u use parenthesis to make it clearer?

OpenStudy (anonymous):

i dont know what parenthesis are

OpenStudy (anonymous):

that's those grouping symbols: ( )

OpenStudy (anonymous):

you know.... when you do a happy face ---> :)

OpenStudy (anonymous):

(|n-1/3n|)(|-1/3|) its the same type of question as before

OpenStudy (anonymous):

sad face :(

OpenStudy (anonymous):

i'm sorry... i gotta go now..... maybe one of these guys/gals can help out????

OpenStudy (anonymous):

pretty please?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

bye bye.... grouping symbol: :)